OpenStudy (anonymous):

(i). F(1) = 1 (ii). F(1) + 2F(2) + 3F(3) + ... + nF(n) = n(n + 1)F(n) for n > 2 calculate the value of F(2015)

OpenStudy (anonymous):

2015(2015+1)F(2015)

OpenStudy (anonymous):

that's just the RHS, what about with the LHS ?

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

F(1) + 2F(2) + 3F(3) + ... + 2015F(2015) = 2015(2015 + 1)F(2015)

OpenStudy (rational):

we have \[ F(1) + 2F(2) + 3F(3) + \cdots + (n-1)F(n-1) + \color{blue}{nF(n)} = n(n + 1)F(n)\] send that blue term to other side and get \[ F(1) + 2F(2) + 3F(3) + \cdots + (n-1)F(n-1) = [n(n + 1) - n]F(n)\] \[ F(1) + 2F(2) + 3F(3) + \cdots + (n-1)F(n-1) = n^2F(n)\] yes ?

OpenStudy (rational):

Next observe that left hand side becomes \(n(n-1)F(n-1)\) and you get a recurrence relation : \[n(n-1)F(n-1) = n^2F(n)\]

OpenStudy (rational):

cancel \(n\) both sides and use induction to conclude \[F(n) = \dfrac{1}{n}\]

OpenStudy (anonymous):

for n = 2014 F(1)+2F(2)+3F(3)+...+2014F(2014)=2014(2014+1)F(2014)=2014(2015)f(2014) for n = 2015 F(1)+2F(2)+3F(3)+...+2015F(2015)=2015(2015+1)F(2015)=2015(2016)f(2015) substract both equation gives us : 2015f(2015) = 2015(2016)f(2015) - 2014(2015)f(2014) still there is the f(2014) --"

OpenStudy (anonymous):

ok, im following till this part n(n - 1)F(n-1) = n^2F(n) (n - 1)F(n-1) = n F(n) but how you establish that F(n) = 1/n ?

OpenStudy (anonymous):

(n - 1)F(n-1) = n F(n) F(n)/F(n - 1) = (n - 1)/n F(n)/F(n - 1) = 1 - 1/n seems the telescopic then, right ?

OpenStudy (anonymous):

F(n)/F(n - 1) = 1 - 1/n n=2 F(2)/F(1) = 1 - 1/2 = 1/2 n=3 F(3)/F(2) = 1 - 1/3 = 2/3 n=4 F(4)/F(3) = 1 - 1/4 = 3/4 . . n=2014 F(2014)/F(2013) = 1 - 1/2014 = 2013/2014 n=2015 F(2015)/F(2014) = 1 - 1/2015 = 2014/2015 multiply all get F(2)/F(1) x F(3)/F(2) x F(4)/F(3) ... F(2014)/F(2013) x F(2015)/F(2014 = 1/2 x 2/3 x 3/4 x ... x 2013/2014 x 2014/2015 F(2015)/F(1) = 1/2015 F(2015) = 1/2015 hehe